1) Using the quadratic equation x2 – 4x – 5 = 0, perform the following tasks:
a) Solve by factoring.
Answer:
Show work in this space.


b) Solve by using the quadratic formula.
Answer:
Show work in this space




Answer by checkley75(3666)

X^2-4X-5=0
(X-5)(X+1)=0
X-5=0
X=5 ANSWER
X+1=0
X=-1 ANSWER
----------------------------------------------
X=(4+-SQRT[-4^2-4*1*-5])/2*1
X=(4+-SQRTR[16+20])/2
X=(4+-SQRT36)/2
X=(4+-6)/2
X=(4+6)/2
X=5 ANSWER
X=(4-6)/2
X=-2/2
X=-1 ANSWER